# Orion: A Fully Homomorphic Encryption Framework for Deep Learning

### v2

@@ 2.5. CKKS Operations
We now describe the primitive CKKS operations: addition, multiplication, rotation, and bootstrapping. CKKS performs modular arithmetic on plaintext and ciphertext polynomials to implement these operations. For fast polynomial multiplication, the Number Theoretic Transformation (NTT) is used to convert the polynomials from the coefficient to the evaluation representation to reduce the running time from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N\log N)$. For this reason, it is common to keep all plaintexts and ciphertexts in the evaluation representation unless otherwise needed. For this section, we will assume we have two cleartext vectors $\mathbf{a},\mathbf{b}\in\mathbb{R}^{n}$ that we encode (as plaintexts $[\mathbf{a}]$ and $[\mathbf{b}]$) both with a scaling factor $\Delta$ and level $\ell$, and further encrypt (as ciphertexts $[\![\mathbf{a}]\!]$ and $[\![\mathbf{b}]\!]$).
@@ 2.5.1. Addition
CKKS supports point-wise addition between one plaintext and one ciphertext ($\mathsf{PAdd}\left([\mathbf{a}],[\![\mathbf{b}]\!]\right)$) or between two ciphertexts (($\mathsf{HAdd}\left([\![\mathbf{a}]\!],[\![\mathbf{b}]\!]\right)$). Each operand must be at the same number of levels and have the same scaling factor. In both $\mathsf{PAdd}$ and $\mathsf{HAdd}$, the output is a ciphertext $[\![\mathbf{c}]\!]_{\text{add}}$ with the same level and scale as the input operands. This ciphertext corresponds to the SIMD addition of the underlying cleartexts: $\mathsf{decode}(\mathsf{decrypt}([\![\mathbf{c}]\!]_{\text{add}}))\approx\mathbf{a}+\mathbf{b}$.
@@ 2.5.2. Multiplication
Similar to addition, CKKS supports point-wise multiplication between one plaintext and one ciphertext ($\mathsf{PMult}\left([\mathbf{a}],[\![\mathbf{b}]\!]\right)$) or between two ciphertexts ($\mathsf{HMult}\left([\![\mathbf{a}]\!],[\![\mathbf{b}]\!]\right)$). Each operand must be at the same number of levels, but not necessarily the same scaling factor. Both $\mathsf{PMult}$ and $\mathsf{HMult}$ result in a ciphertext $[\![\mathbf{c}]\!]_{\text{mult}}$ with a scaling factor $\Delta^{2}$. To prevent exponential growth of the scaling factor, CKKS supports a rescaling procedure that divides the scaling factor by the last prime limb, $q_{\ell}$, and by choosing each $q_{i}\approx\Delta$, the scaling factor remains roughly consistent throughout the computation. Rescaling also removes the last prime limb from each polynomial within $[\![\mathbf{c}]\!]_{\text{mult}}$, effectively reducing the level of the output ciphertext by one.
In addition to rescaling, $\mathsf{HMult}$ requires an expensive key-switching operation (that involves auxiliary evaluation keys) to ensure correct decryption. Key-switching itself is a computationally expensive procedure involving many NTTs and RNS basis conversions [36]. After rescaling (and key-switching for $\mathsf{HMult}$), the output is a ciphertext $[\![\mathbf{c}]\!]_{\text{mult}}$ at level $\ell-1$ and a scaling factor of $\Delta^{2}/q_{\ell}\approx\Delta$. This ciphertext corresponds to the SIMD multiplication of the underlying cleartexts: $\mathsf{decode}(\mathsf{decrypt}([\![\mathbf{c}]\!]_{\text{mult}}))\approx\mathbf{a}\cdot\mathbf{b}$.
@@ 2.5.3. Rotation
Besides addition and multiplication, CKKS also supports the cyclic rotation ($\mathsf{HRot}_{k}\left([\![\mathbf{a}]\!]\right)$) of the underlying cleartext slots by an amount $0<k<n$. Similar to $\mathsf{HMult}$, $\mathsf{HRot}$ also requires a key-switching step using rotation keys so that decryption of the rotated ciphertext is correct. The resulting ciphertext $[\![\mathbf{c}]\!]_{\text{rot}}$ has the same level and scale as the input ciphertext and corresponds to the underlying cleartext vector cyclically rotated ”up” by $k$ slots: $\mathsf{decode}(\mathsf{decrypt}([\![\mathbf{c}]\!]_{\text{rot}})))\approx(a_{k},a_{k+1},\ldots,a_{k-2},a_{k-1})$.
@@ 2.5.4. Bootstrapping
Finally, for a scheme to be fully homomorphic, it must include a way of increasing the number of remaining levels; the bootstrap operation provides this functionality. Bootstrapping is a computationally demanding procedure that increases levels but also consumes a fixed number ($L_{\text{boot}}$) of levels in the process. A ciphertext that begins at level $L$ can only reach an effective level, $L_{\text{eff}}=L-L_{\text{boot}}$ after bootstrapping. A typical $L_{\text{boot}}$ is 15 levels (DHBSGS, ; lattigo, ).
@@ 3. FHE Matrix-Vector Products
In this section, we introduce how efficient homomorphic matrix-vector products are performed within Orion. We focus on the diagonal method of encoding matrices, and its state-of-the-art optimizations (DHBSGS, ) which add both the baby-step giant-step (BSGS) and double-hoisting algorithms. We use these optimizations in every linear transformation in Orion including all linear and convolutional layers of neural networks and the matrix-vector products within bootstrapping. Similar to prior work, and in alignment with our threat model, we assume that the matrix is a cleartext that can be pre-processed before being encoded, and that the vector is a ciphertext.
@@ 3.1. Diagonal Encoding Method
The diagonal method (halevishoup, ) is a straightforward technique for performing homomorphic matrix-vector products. This method works by first extracting the generalized diagonals of a matrix $\mathbf{M}$, defined as $\mathsf{diag}_{\hskip 0.70004ptk}=\mathbf{M}_{[0,\hskip 0.70004ptk]},$ $\mathbf{M}_{[1,\hskip 0.70004ptk+1]},$ $\ldots$,… , $\mathbf{M}_{[w-1,\hskip 0.70004ptk+w-1]}$, where $w$ is the matrix width and the second index is taken modulo $w$. Each diagonal $\mathsf{diag}_{\hskip 0.70004ptk}$ is then multiplied by the input ciphertext, rotated up by $k$ slots, to produce each partial product. Summing all partial products gives the final result. Figure 1(a) illustrates the diagonal method for a $6\times 6$ matrix. Notably, this method requires $n$ ciphertext rotations to perform any $n\times n$ matrix-vector product. Thus, its runtime complexity can be seen as $\mathcal{O}(n)$.
@@ 3.2. Baby-Step Giant-Step Optimization
The baby-step giant-step (BSGS) algorithm reduces the runtime complexity of homomorphic matrix-vector products to $\mathcal{O}(\sqrt{n})$ by leveraging the fact that matrix diagonals can be cheaply rotated before being encoded. More specifically, this algorithm decomposes a matrix-vector product into $n_{1}$ groups of $n_{2}$ diagonals. Each group shares a common (baby-step) ciphertext rotation, and each diagonal within a group is rotated by a different multiple of $n_{1}$. The standard diagonal method is then applied across groups to produce $n_{2}$ partial products, each offset by a multiple of $n_{1}$. Thus, $n_{2}$ (giant-step) ciphertext rotations are required to realign partial products before they are summed to produce the desired result. Equation (1) defines the BSGS algorithm, where $\widetilde{\mathsf{diag}_{k}}=\mathsf{Rot}_{-j\cdot n_{1}}(\mathsf{diag}_{k})$, the $k$’th generalized diagonal of an $n\times n$ matrix, cyclically rotated down by $j\cdot n_{1}$ slots. Figure 1(b) also visualizes the BSGS algorithm when $n_{1}=3$, $n_{2}=2$, and $n_{1}n_{2}=6$, where diagonals are shaded by group. Notably, one can choose any $n_{1}$, $n_{2}$ such that $n_{1}n_{2}=n$, however the number of ciphertext rotations is minimized when $n_{1}=n_{2}=\sqrt{n}$.
@@
Figure 1. Visualizing how the BSGS algorithm (halevishoup, ) reduces the number of ciphertext rotations in matrix-vector products.
@@ 3.3. Hoisting Optimizations
At a high level, hoisting is a cryptographic technique that amortizes the most expensive aspects of the key-switch procedure across many ciphertext rotations. This optimization is only possible when rotating the same ciphertext by different amounts and can therefore only be applied to baby-step rotations in the BSGS algorithm. Bossuat et al. (DHBSGS, ) separately extend this single-hoisting technique to giant-steps and aptly name its implementation double-hoisting. In Orion, every matrix-vector product uses the double-hoisting BSGS algorithm (see Algorithm 6 in (DHBSGS, )). We refer readers to (DHBSGS, ; osiris, ; gazelle, ) for a more detailed explanation of hoisting.
Takeaway: The double-hoisting BSGS method has two fundamental benefits: (1) every matrix-vector product consumes just one multiplicative level, and (2) its performance savings increase with increasing matrix size since BSGS decreases the number of ciphertext rotations from $\mathcal{O}(n)$ to $\mathcal{O}(\sqrt{n})$.
@@ 4. Efficient Convolutions in Orion
In this section, we show how Orion efficiently expresses all convolutions as matrix-vector products to leverage the savings from BSGS and hoisting discussed in Section 3. To achieve this, we introduce a new packing technique called single-shot multiplexing. This technique utilizes a modified Toeplitz formulation of convolutions and is highly effective: it supports arbitrary parameters and halves the level consumption of the multiplexed packing approach of Lee et al. (lee2022, ) while also significantly reducing rotation counts.
Notation: A convolution operation processes a three dimensional input image of size $h_{i}\times w_{i}\times c_{i}$, representing the height, width, and number of channels of the input image, respectively. Then, $c_{o}$ sets of trainable filters of size $f_{h}\times f_{w}\times c_{i}$ are convolved with the input image to produce an output image of size $h_{o}\times w_{o}\times c_{o}$, with parameters such as stride ($s$) and padding ($p$) controlling the output’s dimensions.
Since CKKS operates on one-dimensional vectors of complex (or real) numbers, it is common to first flatten a three-dimensional input into a row-major (raster-scanned) vector of length $h_{i}\hskip 0.50003ptw_{i}\hskip 0.50003ptc_{i}$. For simplicity, we assume that the number of slots, $n$, matches this length so that all data fits perfectly into a single ciphertext. In Section 4.3, we relax this assumption to handle inputs that can be larger than $n$.
@@ 4.1. SISO Convolutions in Orion
We begin with single-input, single-output (SISO) convolutions, where $c_{i}=c_{o}=s=1$, and restrict our attention to same-style convolutions, where the input and output spatial dimensions are equal. An illustrative example is shown in Figure 2. Gazelle (gazelle, ) first introduced a packed SISO method to homomorphically evaluate this convolution, and subsequent works (fhelipe, ; lee2022, ; dacapo, ) have continued to follow this blueprint (Figure 2b). Here, the input is flattened, packed into a ciphertext, and cyclically rotated $f_{h}\hskip 0.50003ptf_{w}$ times. Each rotated ciphertext is then multiplied by a punctured plaintext, which encodes $n$ copies of a unique filter weight and is selectively zeroed to ensure only the pixels that interact with the filter weight are processed. Finally, the partial products from each multiplication are summed to produce the final output.
@@
Orion: Our key observation is that prior works’ packed SISO method is equivalent to the diagonal method: ciphertexts are rotated and multiplied with pre-processed plaintexts. We show this by working backwards from Figure 2b to derive its analogous matrix-vector product. In doing so, we arrive at the Toeplitz formulation of a convolution, shown in Figure 2a. In this representation, the filter expands into a weight matrix of size $h_{o}\hskip 0.50003ptw_{o}\times h_{i}\hskip 0.50003ptw_{i}$, where each row corresponds to one filter multiplication. For example, the first row of the matrix in Figure 2a performs a dot product between the filter weights {5,6,8,9} and the input pixels {a,b,d,e} which corresponds to the filter being placed at the top left corner of the input with padding $p=1$. Subsequent rows of the Toeplitz matrix are generated as the kernel slides across the input image. Performing the diagonal encoding method on the matrix in Figure 2a produces the exact operations seen in the packed SISO method in Figure 2b.
The fundamental benefit of this observation is that we can build a Toeplitz matrix for any arbitrary convolution and apply the double-hoisting BSGS algorithm when evaluating its matrix-vector product. To the best of our knowledge slytHErin (slytherin, ) is the only prior work which proposes a similar strategy. With this, Orion reduces the number of ciphertext rotations for any SISO convolution from $\mathcal{O}(f)$ to $\mathcal{O}(\sqrt{f})$, where $f$ is the total number of filter elements.

### v3

@@ 2.5. CKKS Operations
We now describe the primitive CKKS operations: addition, multiplication, rotation, and then bootstrapping. CKKS performs modular arithmetic on plaintext and ciphertext polynomials to implement these operations. For fast polynomial multiplication, the Number Theoretic Transform (NTT) is used to convert the polynomials from the coefficient to the evaluation representation to reduce the running time from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N\log N)$. For this reason, it is common to keep all plaintexts and ciphertexts in the evaluation representation unless otherwise needed. For this section, we will assume we have two cleartext vectors $\mathbf{a},\mathbf{b}\in\mathbb{R}^{n}$ that we encode (as plaintexts $\mathbf{a}]$ and $\mathbf{b}]$) both with a scaling factor $\Delta$ and level $\ell$, and further encrypt (as ciphertexts $\mathbf{a}]\!]$ and $\mathbf{b}]\!]$).
@@ 2.5.1. Addition
CKKS supports point-wise addition between one plaintext and one ciphertext ($\mathsf{PAdd}\left([\mathbf{a}],[\![\mathbf{b}]\!]\right)$) or between two ciphertexts $($$\mathsf{HAdd}\left([\![\mathbf{a}]\!],[\![\mathbf{b}]\!]\right)$). Each operand must be at the same number of levels and have the same scaling factor. In both $\mathsf{PAdd}$ and $\mathsf{HAdd}$, the output is a ciphertext $\mathbf{c}]\!]_{\text{add}}$ with the same level and scale as the input operands. This ciphertext corresponds to the SIMD addition of the underlying cleartexts: $\mathsf{decode}(\mathsf{decrypt}([\![\mathbf{c}]\!]_{\text{add}}))\approx\mathbf{a}\oplus\mathbf{b}$.
@@ 2.5.2. Multiplication
Similar to addition, CKKS supports point-wise multiplication between one plaintext and one ciphertext ($\mathsf{PMult}\left([\mathbf{a}],[\![\mathbf{b}]\!]\right)$) or between two ciphertexts ($\mathsf{HMult}\left([\![\mathbf{a}]\!],[\![\mathbf{b}]\!]\right)$). Each operand must be at the same number of levels, but not necessarily the same scaling factor. Both $\mathsf{PMult}$ and $\mathsf{HMult}$ result in a ciphertext $\mathbf{c}]\!]_{\text{mult}}$ with a scaling factor $\Delta^{2}$. To prevent exponential growth of the scaling factor, CKKS supports a rescaling procedure that divides the scaling factor by the last prime limb, $q_{\ell}$, and by choosing each $q_{i}\approx\Delta$, the scaling factor remains roughly consistent throughout the computation. A key property of the rescaling procedure is that for each coefficient, the component belonging to $\mathbb{Z}_{q_{\ell}}$ can be (and is) discarded, so that the resulting polynomial has coefficients in $\mathbb{Z}_{Q_{\ell-1}}$ rather than $\mathbb{Z}_{Q_{\ell}}$; this is why multiplication is said to consume a level.
In addition to rescaling, $\mathsf{HMult}$ requires an expensive key-switching operation (that involves auxiliary evaluation keys) to ensure correct decryption. Key-switching itself is a computationally expensive procedure involving many NTTs and RNS basis conversions (Han and Ki, 2020; Neda et al., 2024). After rescaling (and key-switching for $\mathsf{HMult}$), the output is a ciphertext $\mathbf{c}]\!]_{\text{mult}}$ at level $\ell-1$ and a scaling factor of $\Delta^{2}/q_{\ell}\approx\Delta$. This ciphertext corresponds to the SIMD multiplication of the underlying cleartexts: $\mathsf{decode}(\mathsf{decrypt}([\![\mathbf{c}]\!]_{\text{mult}}))\approx\mathbf{a}\odot\mathbf{b}$.
@@ 2.5.3. Rotation
Besides addition and multiplication, CKKS also supports the cyclic rotation ($\mathsf{HRot}_{k}\left([\![\mathbf{a}]\!]\right)$) of the underlying cleartext slots by an amount $0<k<n$. Similar to $\mathsf{HMult}$, $\mathsf{HRot}$ also requires a key-switching step using rotation keys so that decryption of the rotated ciphertext is correct. The resulting ciphertext $\mathbf{c}]\!]_{\text{rot}}$ has the same level and scale as the input ciphertext and corresponds to the underlying cleartext vector cyclically rotated ”up” by $k$ slots: $\mathsf{decode}(\mathsf{decrypt}([\![\mathbf{c}]\!]_{\text{rot}})))\approx(a_{k},a_{k+1},\ldots,a_{k-2},a_{k-1})$.
@@ 2.5.4. Bootstrapping
Finally, for a scheme to be fully homomorphic, it must include a way of increasing the number of remaining levels; the bootstrap operation provides this functionality. Bootstrapping is a computationally demanding procedure that increases levels but also consumes a fixed number ($L_{\text{boot}}$) of levels in the process. Therefore, a ciphertext that begins at level $L$ can only reach an effective level, $L_{\text{eff}}=L-L_{\text{boot}}$ after bootstrapping. A typical $L_{\text{boot}}$ is between 13 to 15 levels (Bossuat et al., 2020; lat, 2022; HEA, 2016).
@@ 3. FHE Matrix-Vector Products
In this section, we introduce how efficient homomorphic matrix-vector products are performed within Orion. We focus on the diagonal method (Halevi and Shoup, 2014) of encoding matrices and its state-of-the-art optimizations that include both the baby-step giant-step (BSGS) (Halevi and Shoup, 2014) and double-hoisting algorithms (Bossuat et al., 2020). These are powerful algorithmic optimizations that significantly reduce compute costs. We use these optimizations in every linear transformation in Orion including all fully-connected and convolutional layers of neural networks and the matrix-vector products within bootstrapping. Similar to prior work, and in alignment with our threat model, we assume that the weight matrix is a cleartext that can be pre-processed before being encoded, and that the input vector is a ciphertext.
@@ 3.1. Diagonal Encoding Method
The diagonal method (Halevi and Shoup, 2014) is one technique for performing homomorphic matrix-vector products. This method works by first extracting the generalized diagonals of a matrix $\mathbf{M}$, defined as $\mathsf{diag}_{\hskip 0.81949ptk}=\mathbf{M}_{[0,\hskip 0.81949ptk]},$ $\mathbf{M}_{[1,\hskip 0.81949ptk+1]},$ $\ldots,$ $\mathbf{M}_{[w-1,\hskip 0.81949ptk+w-1]}$, where $w$ is the matrix width and the second index is taken modulo $w$. Each diagonal, $\mathsf{diag}_{\hskip 0.81949ptk}$, is then multiplied by the input ciphertext, rotated up by $k$ slots, to produce each partial product. Summing all partial products gives the final result. Figure 2(a) illustrates the diagonal method for a $6\times 6$ matrix. Notably, this method requires $n$ ciphertext rotations to perform any $n\times n$ matrix-vector product. Thus, its runtime complexity can be seen as $\mathcal{O}(n)$.
Figure 2. Visualizing how the BSGS algorithm (Halevi and Shoup, 2014) reduces the number of ciphertext rotations in matrix-vector products.
@@ 3.2. Baby-Step Giant-Step Optimization
The baby-step giant-step (BSGS) algorithm reduces this complexity to $\mathcal{O}(\sqrt{n})$ by leveraging the fact that matrix diagonals can be cheaply rotated before being encoded. More specifically, BSGS decomposes a matrix-vector product into $n_{1}$ groups of $n_{2}$ diagonals. Each group shares a common (baby-step) ciphertext rotation, and each diagonal within a group is rotated by a different multiple of $n_{1}$. The standard diagonal method is then applied across groups to produce $n_{2}$ partial products, each offset by a multiple of $n_{1}$. Thus, $n_{2}$ (giant-step) ciphertext rotations are required to align partial products before they are summed to produce the desired result. Equation (1) defines the BSGS algorithm, where $\widetilde{\mathsf{diag}_{k}}=\mathsf{Rot}_{-\hskip-0.81949ptj\cdot n_{1}}(\mathsf{diag}_{k})$, the $k$’th generalized diagonal of an $n\times n$ matrix, cyclically rotated down by $j\cdot n_{1}$ slots. Figure 2(b) also visualizes the BSGS algorithm when $n_{1}=3$, $n_{2}=2$, and $n_{1}n_{2}=6$, where diagonals are shaded by group. Notably, one can choose any $n_{1}$, $n_{2}$ such that $n_{1}n_{2}=n$, however the number of ciphertext rotations is minimized when $n_{1}=n_{2}=\sqrt{n}$.
@@ 3.3. Hoisting Optimizations
At a high level, hoisting is a cryptographic technique that reuses the most expensive aspects of the key-switch procedure over many ciphertext rotations. This optimization is only possible when rotating the same ciphertext by different amounts and can therefore only be applied to baby-step rotations in the BSGS algorithm. Bossuat et al. (Bossuat et al., 2020) separately extend this single-hoisting technique to giant-steps and aptly name its implementation double-hoisting. In Orion, every matrix-vector product uses the double-hoisting BSGS algorithm (see Algorithm 6 in (Bossuat et al., 2020)). We refer readers to (Bossuat et al., 2020; Ebel and Reagen, 2024; Juvekar et al., 2018) for a more detailed explanation of hoisting.
Takeaway: The double-hoisting BSGS method has two fundamental benefits: (i) every matrix-vector product consumes just one multiplicative level, and (ii) its performance savings increase with increasing matrix size since BSGS decreases the number of ciphertext rotations from $\mathcal{O}(n)$ to $\mathcal{O}(\sqrt{n})$.
@@ 4. Efficient Convolutions in Orion
In this section, we show how Orion efficiently expresses all convolutions as matrix-vector products to leverage the savings from BSGS and hoisting discussed in Section 3. To achieve this, we introduce a new packing technique called single-shot multiplexing. This technique utilizes a modified Toeplitz formulation of convolutions and is highly effective: it supports arbitrary parameters and halves the level consumption of strided convolutions compared to the multiplexed packing approach of the state-of-the-art (Lee et al., 2022b) while also significantly reducing rotation counts.
Since CKKS operates on vectors of complex (or real) numbers, it is common to first flatten a three-dimensional convolutional input into a row-major (raster-scanned) vector of length $h_{i}\hskip 0.50003ptw_{i}\hskip 0.50003ptc_{i}$, following the notation from Section 2. For simplicity, we assume that the number of slots, $n$, matches this length so that all data fits perfectly into a single ciphertext. In Section 4.3, we relax this assumption to handle inputs larger than $n$.
@@ 4.1. SISO Convolutions in Orion
We begin with single-input, single-output (SISO) convolutions where $c_{i}=c_{o}=s=1$ and restrict our attention to same-style convolutions, where the input and output spatial dimensions are equal. An illustrative example is shown in Figure 3. Gazelle (Juvekar et al., 2018) first introduced a packed SISO method to homomorphically evaluate this convolution, and subsequent works (Krastev et al., 2024; Lee et al., 2022b; Cheon et al., 2024) have continued to follow this blueprint (Figure 3b). Here, the input is flattened, packed into a ciphertext, and cyclically rotated $f_{h}\hskip 0.50003ptf_{w}$ times. Each rotated ciphertext is then multiplied by a punctured plaintext, which encodes $n$ copies of a unique filter weight and is selectively zeroed to ensure only the pixels that interact with the filter weight are processed. Finally, the partial products from each multiplication are summed to produce the final output.
Orion: Our observation here is that prior works’ packed SISO method is equivalent to the diagonal method: ciphertexts are rotated and multiplied with pre-processed plaintexts. We show this by working backwards from Figure 3b to derive its analogous matrix-vector product. In doing so, we arrive at the Toeplitz formulation of a convolution, shown in Figure 3a. In this representation, the filter expands into a weight matrix of size $h_{o}\hskip 0.50003ptw_{o}\times h_{i}\hskip 0.50003ptw_{i}$, where each row corresponds to one filter multiplication. For example, the first row of the matrix in Figure 3a performs a dot product between the filter weights {5,6,8,9} and the input pixels {a,b,d,e}, which corresponds to the filter being placed at the top left corner of the input with padding $p=1$. Subsequent rows of the Toeplitz matrix are generated as the kernel slides across the input image. Performing the diagonal encoding method on the matrix in Figure 3a produces the exact operations seen in the packed SISO method in Figure 3b.
@@
The fundamental benefit of this observation is that we can build a Toeplitz matrix for any arbitrary convolution and then apply the double-hoisting BSGS algorithm when evaluating its matrix-vector product. To the best of our knowledge slytHErin (Intoci et al., 2023) is the only prior work that proposes a similar strategy. With this, Orion reduces the number of ciphertext rotations for any SISO convolution from $\mathcal{O}(f)$ to $\mathcal{O}(\sqrt{f})$, where $f$ is the total number of filter elements.